could someone please help me divide -15/-3
Answer by jim_thompson5910(35256)

since 3*5 = 15, this means that 15/3 = 5


So -15/(-3) = 5


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim